OpenStudy (anonymous):

The main tank at an aquaruim is a cylinder with a diameter of 203 ft and a height of 25 ft. Find the volume of the tank to the nearest cubic foot. Use the pi key on your calculator for the value pi... A) 794,234 ft^3 ... B) 809,137 ft^3 ... C) 815,297 ft^3 ... D) 799,247 ft^3

OpenStudy (nubeer):

Volume of Cylinder =pi*r^2*h r=dia/2 you have all the values just plug in and get the answer.

OpenStudy (anonymous):